OpenStudy (anonymous):

What is the simplest form of the product? 3^ sqrt 4x^2 * 3^ sqrt 8x^7 2x^3 * 3^ sqrt 4 3^ sqrt 32x^9 2x^3 * 3^ sqrt 4x^9 none of these

OpenStudy (anonymous):

@amistre64 ; help please ? (:

OpenStudy (amistre64):

can you place ( ) around the stuff that is inside the radicals? trying to read this is a pain

OpenStudy (amistre64):

does 3^ sqrt mean cube root?

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

sorry . . . .

OpenStudy (amistre64):

oy vey!! its simpler if you short it as cbrt(..), or even 3rt(..)

OpenStudy (amistre64):

3rt(4x^2) * 3rt(8x^7) is this the problem?

OpenStudy (anonymous):

sorry , i didnt know i could write it like that. lol

OpenStudy (anonymous):

& yes thats the problem (:

OpenStudy (amistre64):

hmm, it might be prudent to use an exponential form for the radicals;\[(4x^2)^{1/3} * (8x^7)^{1/3}\] the product of an exponented thing is equal to the product of the separate terms exponented \[4^{1/3}*(x^2)^{1/3} * 8^{1/3}*(x^7)^{1/3}\] another property is that the exponent of and exponent is just the product of the exponents \[4^{1/3}*x^{2/3} * 8^{1/3}*x^{7/3}\] and rearrange \[4^{1/3} * 8^{1/3}*x^{7/3}*x^{2/3}\] \[\large(4*8)^{1/3}*x^{7/3+2/3}\]

OpenStudy (amistre64):

\[\large(4*8)^{1/3}*x^{7/3+2/3}\] \[\large(32)^{1/3}*x^{(7+2)/3}\] \[\large\sqrt[3]{32}*x^{9/3}\] \[\large\sqrt[3]{32}*(x^{9})^{1/3}\] \[\large\sqrt[3]{32}*\sqrt[3]{x^{9}}\] \[\large\sqrt[3]{32x^{9}}\]

OpenStudy (amistre64):

now, if thats "simplest form" or not i cant determine, but its at lesast one of the answer options

OpenStudy (anonymous):

ohhh , I see . Thank you sooo much for helping me . i have been stuck on this for sooo long . . . lol

OpenStudy (amistre64):

\[\large\sqrt[3]{32x^{9}}=\large x^3\sqrt[3]{32}\]so its up to you what "simplest" form is :)

OpenStudy (amistre64):

hmm, and 32 = 8*4 and 8=2^3 hmmm

OpenStudy (amistre64):

\[\large\sqrt[3]{32x^{9}}=\large x^3\sqrt[3]{2^3*4}=2x^3\sqrt[3]{4}\]that should do it

OpenStudy (amistre64):

gotta memorize the properties, or at least understand the underlying math that makes them work in order to recreate them ;) good luck